OpenStudy (emiiamontalvo420):

Write the equation of a line which contains the points (2,2) and (-3,12) Remember to first find the slope and then use the point slope formula y - y1 = m(x - x1)

OpenStudy (royalranger):

\[m = \frac{y _{2} - y _{1} }{ x _{2}- x _{1} }\]

OpenStudy (sunnnystrong):

as @royalranger said... The equation for finding the slope of a line when given two points is \[m=(y2-y1)\div(x2-x1) \] (2,2) and (-3,12) are our points. Just plug the two points into the equation.... \[m=(2-12)/(2+3) \] Once you solve for m. y - y1 = m(x - x1) Use one of your points and your slope to find the equation of the line.